OpenStudy (alfred565):

Help??? I left my book at school V varies inversely with T and V = 18 when T = 3. Which equation shows this relationship?

OpenStudy (anonymous):

inverse variation: y = k / x V = k / T substitute and solve for k- the constant of variation, 18 = k / 3 k = 54 the equation is: V = 54/T

OpenStudy (anonymous):

k= 54/T as she said